# Integrating Biological and Machine Intelligence: Attention Mechanisms in Brain-Computer Interfaces

### v1

@@ 2.1 Types of Attention Modules
(1) Channel Attention Module. The channel attention module is designed to assess and adaptively weight the importance of each EEG channel. Given that different brain regions contribute unequally to various tasks, the channel attention module enables the model to prioritize channels with the most relevant information while minimizing the impact of less informative channels, thereby improving task-specific analysis and reducing noise.
For a given multi-channel EEG signal vector $X_{c}\in\mathbb{R}^{1\times C}$, where $C$ denotes the number of channels, an attention weight vector $W_{c}\in\mathbb{R}^{1\times C}$ is initialized. The model calculates a weighted combination of $X_{c}$ and $W_{c}$, followed by the application of the softmax function to produce the output $\hat{X}_{c}$ from the channel attention module, as shown below:
$\hat{X}_{c}=\text{softmax}(X_{c}\odot W_{c})$ (1)
In BCI applications, the channel attention module effectively identifies the most relevant brain regions for specific tasks, such as motor imagery or emotional state classification [25, 26, 27, 28]. By concentrating on channels linked to critical brain functions, this module improves feature extraction, leading to enhanced performance in BCI systems and other brain modeling applications.
(2) Temporal Attention Module. The temporal attention module is designed to effectively capture the dynamic fluctuations in brain activity that occur over time, acknowledging that EEG signals reflect varying brain states depending on the nature of a given task. During task execution, brain activity often exhibits multiple phases of relevance, with some fluctuations directly related to the task while others contribute less meaningful information. By assigning higher attention weights to time periods that are directly correlated with the task at hand, the module enhances the model’s ability to extract task-relevant features and reduce the influence of irrelevant temporal variations.
For a given EEG signal vector $X_{t}\in\mathbb{R}^{1\times T}$, where $T$ represents the temporal dimension, an attention weight vector $W_{t}\in\mathbb{R}^{1\times T}$ is initialized. The model then calculates a weighted combination of $X_{t}$ and $W_{t}$, followed by the application of the softmax function to yield the temporal attention output $\hat{X}_{t}$, as shown below:
$\hat{X}_{t}=\text{softmax}(X_{t}\odot W_{t})$ (2)
In BCI applications, the temporal attention module effectively captures relevant patterns of brain activity during cognitive or emotional tasks [29, 30]. For example, in motor imagery tasks, it highlights critical moments by identifying peaks in attention at significant time intervals [31, 32]. By concentrating on time segments that are closely linked to the task, the temporal attention module ensures that the model accurately captures the dynamic characteristics of EEG signals, leading to improved performance in analyzing time-series brain activity data and enhancing the robustness of BCI systems.
(3) Frequency Attention Module. The frequency attention module is designed to leverage the significance of frequency components in EEG signals, as frequency-domain analysis reveals crucial insights into neural activities across different scales. Features such as Power Spectral Density (PSD) and Differential Entropy (DE) are valuable indicators of brain activity, and the frequency attention module adaptively highlights these key features by assigning weights to various frequency components within the signal.
For a frequency vector $X_{f}\in\mathbb{R}^{1\times F}$, where $F$ represents the frequency dimension, an attention weight vector $W_{f}\in\mathbb{R}^{1\times F}$ is initialized. The model computes a weighted combination of $X_{f}$ and $W_{f}$, and applies the softmax function to derive the output $\hat{X}_{f}$ from the frequency attention module, as shown below:
$\hat{X}_{f}=\text{softmax}(X_{f}\odot W_{f})$ (3)
In BCI applications, the frequency attention module effectively identifies and enhances the most informative frequency components, such as Alpha, Beta, and Gamma bands, which are known to be critical in tasks like attention monitoring, motor imagery, or emotional state classification [33, 34]. By adaptively assigning higher weights to these significant frequency bands, the module ensures that the model focuses on the most relevant aspects of the EEG data, ultimately enhancing the feature extraction process. This selective attention to key spectral features contributes to improved classification accuracy and robustness in BCI systems, thereby advancing the reliability of EEG-based analyses.
@@ 2.2 Embedding Methods of Attention Modules
For example, embedding a channel attention module specifically enhances the model’s ability to capture important features at the channel level. Du et al. proposed the ATDD-LSTM model, which combined a channel attention module with a long short-term memory (LSTM) network. In this approach, the channel attention module was applied to feature vectors extracted by the LSTM layers, allowing the model to concentrate on channels most relevant to specific emotions while downplaying less relevant ones, which improved the accuracy of emotion recognition [35]. Inspired from this, Xu et al. integrated channel attention into a graph convolutional network (GCN), taking into account the spatial relationships between EEG recording electrodes [36]. Beyond channel attention, some studies focus on embedding temporal attention module to emphasize the temporal dynamics of EEG signals. Zhang et al. proposed a convolutional recurrent attention model that used CNNs to encode high-level representations and combined them with recurrent attention mechanisms (including LSTM networks and temporal attention modules). This method calculated attention weights on dynamic temporal features, allowed the model to focus on the most informative time periods, and extracted more valuable temporal information [37]. Inspired by the psychological peak-end rule, Kim et al. developed a model that integrated a bidirectional LSTM network with a temporal attention module. It assigned greater weight to emotional states occurring at key moments, capturing the dynamic variability of emotions over time and enhancing the model’s interpretability [38]. For frequency-domain features, frequency attention is rarely modeled in isolation. Instead, it is typically integrated with spatial and temporal features to enhance EEG signal representation.
@@
For example, Tao et al. proposed a deep learning model that incorporated both channel attention and inter-sample attention mechanisms. Since the samples were segmented based on time, the inter-sample attention effectively functioned as temporal attention. This approach enabled the model to effectively prioritize significant information across different channels and temporal segments for feature extraction [27]. Extending beyond channel and temporal information, Jia et al. introduced a spatio-temporal-spectral attention dense network that simultaneously considered temporal, frequency, and spatial features. This model adaptively captured crucial information across brain regions (spatial, i.e., channels), frequency bands (spectral), and time, providing a comprehensive feature extraction framework [30]. Xiao et al. extended Jia et al.’s model by proposing a neural network based on 4D attention [34]. In this approach, the channel dimension of the input samples is transformed into a two-dimensional feature to preserve the spatial positional information of the EEG signal electrodes, while also incorporating the time and frequency dimensions. It computed spatial attention (addressing the spatial positional relationships between channels) and frequency attention (applied to power spectral density and differential entropy features). These attention weights were then applied to refine the input, resulting in enhanced output features. Jia et al. ’s and Xiao et al. ’s models both leveraged temporal, frequency, and spatial characteristics of EEG channels, calculating attention weights across these dimensions to enhance the model’s focus on information pertinent to specific tasks. The main difference between their approaches lies in how they calculated attention weights and structured their models. Jia et al. computed attention separately across the frequency and time dimensions in parallel, and then merged these features for classification. In contrast, Xiao et al. integrated all dimensions into a unified 4D representation before computing attention, providing a different approach to capture feature interdependencies.
@@
Besides of the aforementioned methods, Cai et al. introduced a dynamic attention mechanism that assigned different weights to different frequency sub-bands and channels of EEG signals [33]. This dynamic approach optimized feature representation and was applied within an adaptive decoding framework for complex downstream tasks. A crucial aspect of applying attention mechanisms is defining the shape and dimensions of the input feature matrix. In previous research, attention weight computations have predominantly relied on the strict Euclidean geometric space of the feature matrix. However, given the brain’s complex topological structure, using Euclidean space alone may not accurately capture its underlying properties. Consequently, several studies have sought to align feature matrix definitions more precisely with the brain’s physiological structure by incorporating non-Euclidean space representations within attention mechanisms. For example, Zhang et al. introduced the concept of manifolds, proposing a time-frequency domain feature learning model that integrated both Riemannian manifold and Euclidean space representations. Their work demonstrated the effectiveness of attention mechanisms in synthesizing feature information across different mathematical domains [39]. Additionally, to better capture the spatial relationships among EEG electrode channels, Jia et al. ’s GraphSleepNet [40] and Zhang et al. ’s hierarchical attention network based on graph structures [41] both utilized GCNs to model the spatial relationships of EEG electrodes. These approaches leveraged attention mechanisms across both time and space, significantly enhancing performance in tasks such as sleep stage classification and movement intention prediction.
@@
[27] 2020 Channel,Time RNN emotion recognition DEAP,DREAMER
[37] 2019 Time CNN,LSTM subject-independent movement intention recognition BCI competition IV dataset 2a
[38] 2020 temporal LSTM,CNN emotion recognition DEAP
[30] 2020 Channel,Time,Frequency CNN emotion recognition SEED,SEEDIV
[34] 2022 Channel,Time,Frequency LSTM,CNN emotion recognition DEAP,SEED,SEEDIV
[33] 2021 Frequency,Channel CNN AAD KUL,DTU
[39] 2020 Time,Frequency LSTM,CNN emotion recognition SEED-VIG,SEED, BCI-IV 2A,BCI-IV 2B
[40] 2020 Time,Channel GCN Sleep stage recognition MASS-SS
[41] 2019 Channel,Time CNN,GCN left/right fist open and close intentions PhysioNet dataset

### v2

@@ 2.1 Types of Attention Modules
(1) Channel Attention Module.
The channel attention module is designed to assess and adaptively weight the importance of individual EEG channel. The brain, as a complex biological system, relies on specialized brain regions that fulfill distinct roles while maintaining inter-regional interactions to adapt to diverse real world tasks[25]. Extensive research from neuroscience and computational studies indicates that different brain areas contribute unequally to specific cognitive functions. In BCI applications, channel attention module effectively identifies and emphasizes the most informative brain regions for tasks such as motor imagery, emotion recognition, and visual perception[26, 27, 28, 29, 30, 31, 32]. By assigning weights to EEG channels based on their task relevance, it allows models to focus on salient signals from informative channels and suppress noises from less informative channels. The most straightforward approach to implement the traditional attention mechanism to the channel dimension of electroencephalography (EEG) signals is to initialize trainable weight parameters for each channel. These weights are multiplied with the corresponding EEG input features and dynamically updated during model training, Higher weights signify greater channel importance for the task. For example, Huang et al. introduced a channel attention layer within a CNN-Bi-LSTM architecture to balance contributions of EEG channels in emotion recognition[33]. Su et al. designed a soft attention mechanism that generates channel masks by optimizing task objectives for auditory attention detection in multi-speaker scenarios[34]. Hsu et al. proposed the Deep EEG-Channel-Attention (DEC) module, which adaptively adjusts channel weights in motor imagery classification[35]. Similar efforts by Wimpff et al. and Zhang et al. further validate the utility of channel attention mechanism[36, 37]. Mathematically, given a multi-channel EEG input $X_{c}\in\mathbb{R}^{C\times T}$($C$ denotes the number of channels and $T$ is the number of time points), a trainable attention weight vector $W_{c}\in\mathbb{R}^{1\times C}$ is introduced. The weighted output $\hat{X}_{c}$ is computed as:
$\hat{X}_{c}=\sigma(X_{c}\odot W_{c})$ (1)
where $\sigma$ denotes an activation function (e.g., Sigmoid or Softmax), and $\odot$ denotes element-wise multiplication.
To capture the spatial dependencies between EEG electrodes, channel attention can be integrated with Graph Neural Networks (GNNs) [38]. GNNs are particularly well-suited for EEG data due to their ability to model non-Euclidean structures, such as the irregular spatial layout of scalp electrodes. They represent EEG channels as nodes and their spatial relationships as edges in a graph. For example, Lin et al. demonstrated that retaining only 20% of EEG channels via attention-based selection within a GNN architecture yielded over 90% accuracy on the DEAP and SEED datasets. Even when limited to the top 12 channels, performance declined by only 1.3% on SEED [39]. Additional channel-specific weighting modules can be introduced prior to GNN input, further enhancing flexibility and performance[40]. The general formulation for channel attention in GNNs is as follows:
$H(x)=\sigma(\tilde{\mathbf{A}}F(X\odot W_{c})W_{G})$ (2)
Here, $\sigma$ represents an activation function (e.g. sigmoid or softmax), $\tilde{\mathbf{A}}$ is the adjacency matrix, $F(.)$ is a mapping function (e.g., convolutional or fully connected layer), and $W_{c}$ is a learnable GNN parameter, and the rest are as previously defined.
Recent advances in channel-specific attention mechanisms for EEG have focused on enhancing spatial modeling without relying on temporal or spectral features. By concentrating on channels associated with key brain functions, these mechanisms improve feature extraction and boost the performance of brain-computer interface systems and other neural decoding applications. A notable innovation is the use of dynamic, data-driven channel connectivity, where traditional fixed adjacency matrices based on anatomical electrode positions are replaced by adaptive graph structures learned directly from data. This flexibility allows the model to assign edge weights via attention scores that reflect task-specific functional relevance. For example, it emphasizes frontal-parietal connections during emotional processing, which could capture context-dependent neural couplings unique to each individual or task. Another promising direction involves sparse and interpretable channel weighting through regularization techniques such as L1-norm, which encourage the selection of a minimal subset of critical channels. This reduces redundancy while aligning with established neurophysiological insights, enhancing model transparency. Hierarchical channel grouping provides further refinement by organizing electrodes into brain regions and applying attention at both inter-regional and intra-regional levels, effectively capturing multi-scale spatial dependencies in neural activity. To address inter-subject variability, cross-subject adaptive attention mechanisms incorporate meta-learning or transfer learning strategies to personalize channel weighting for new individuals, leveraging common neural topologies while accommodating unique signal characteristics. Finally, lightweight attention architectures have been developed to ensure computational efficiency in real-time applications by incorporating compact modules such as bottleneck layers that select essential channels with minimal latency. Collectively, these innovations significantly advance the modeling of EEG spatial dynamics, enabling more adaptive, interpretable, and efficient neural systems.
(2) Temporal Attention Module.
The temporal attention module is designed to effectively capture the dynamic fluctuations in brain activity over time, acknowledging that EEG signals reflect temporally varying neural states depending on the nature of the task. During task execution, cognitive processes unfold across multiple phases, with certain moments carrying higher relevance while others contribute less informative or even distracting signals. For example, in emotion recognition tasks, the onset, peak, and offset of emotional responses are distributed across distinct time points, reflecting task-specific temporal patterns[41, 42]. Findings in neuroscience and cognitive science have consistently demonstrated that different time steps in EEG signals contribute unevenly to cognitive function processing[43, 44, 45]. By assigning higher attention weights to task-relevant time intervals and suppressing irrelevant fluctuations, the temporal attention module enhances the model’s capacity to extract discriminative temporal features. It ultimately improves the robustness and precision of brain activity analysis in time-series modeling.
Due to the inherently non-stationary nature of EEG signals, two fundamental challenges arise when applying temporal attention: determining the appropriate duration of each time window and defining the optimal number of segments necessary to isolate meaningful neural responses. These parameters often vary across tasks and recording conditions, making them more complex to optimize than the fixed and predefined spatial structure seen in channel attention mechanisms. To address these issues, researchers have proposed various solutions. Jiang et al. introduced a dynamic weighting strategy for identifying key temporal fragments, demonstrating that task performance, such as emotion recognition, can be significantly influenced by the selected window length[46]. Their work suggested that optimal segment durations should be adapted based on task characteristics, such as distinguishing between structured emotional induction and natural conversation. Yang et al. developed a Gated Temporal-Separable Attention Network that leverages causal and dilated convolutions to extract multi-scale temporal features, subsequently assigning attention weights to relevant segments [47]. Beyond these examples, temporal attention has been widely applied in diverse BCI contexts, including cognitive and emotional state monitoring[48, 49] as well as motor imagery classification [50, 51]. A general representation of temporal attention in this context can be formalized as follows. Given an EEG signal vector $X_{t}\in\mathbb{R}^{1\times T}$ ($T$ is the length of the temporal sequence), an attention weight vector $W_{t}\in\mathbb{R}^{1\times(n\cdot t)}$ is initialized. Here, $n$ denotes the number of time windows, and $t$ is the length of each window, with $n\cdot t=T$. These parameters can either be fixed as hyperparameters or learned dynamically. The weighted output $\hat{X}_{t}$ is computed as:
$\hat{X}_{t}=\sigma(X_{t}\odot W_{t})$ (3)
where $\sigma$ is an activation function, and $\odot$ denotes element-wise multiplication. This formulation enables the model to focus on critical time segments that are aligned with cognitive task demands. Furthermore, research in temporal attention for EEG is increasingly oriented toward dynamic, neuroscience-inspired adaptability and data-efficient unsupervised learning. One promising direction involves drawing from neural coding principles, such as the variable temporal resolution observed in biological spiking neural networks, to enable models to adjust temporal granularity in a task-sensitive manner, eliminating the need for heuristic fixed windows. Additionally, unsupervised learning approaches such as contrastive learning or masked segment prediction offer a pathway for discovering generalizable temporal patterns from large volumes of unlabeled EEG data. These methods help mitigate the limitations of scarce annotated datasets while extracting meaningful representations from latent brain dynamics. To further enhance interpretability, generative models can be used to visualize attention-aligned neural features, and causal inference techniques may uncover how specific time segments contribute to cognitive or behavioral outcomes. Together, these innovations offer a pathway toward more adaptive, efficient, and explainable EEG-based systems, with broad implications for clinical neuroscience and next-generation brain-computer interfaces.
(3) Frequency Attention Module.
The frequency attention mechanism is designed to capture and enhance the spectral characteristics of EEG signals, which reflect brain activity across different cognitive and behavioral states. Frequency-domain analysis has long been recognized as a powerful approach in neuroscience and signal processing. Decades of research in neuroscience and signal processing have shown that specific frequency bands, namely $\delta$ (0.5-4 Hz), $\theta$ (4 - 8 Hz), $\alpha$ (8 - 12 Hz), $\beta$ (13 - 30 Hz), and $\gamma$ ($>$ 30 Hz), are closely associated with distinct brain functions. In practical BCI applications, this predefined division of frequency bands is often used as prior knowledge to guide the analysis of task-related EEG features. The frequency attention module builds on this foundation by learning to adaptively assign higher weights to the most informative spectral components, such as power spectral density (PSD) or differential entropy (DE), thereby improving the extraction of features that are most relevant to tasks like auditory attention detection, motor imagery, and emotional state classification[52, 53].
In practical use, the frequency attention mechanism is typically implemented by first decomposing EEG signals into multiple sub-bands based on prior neuroscience knowledge, and then applying attention weights to each band. For example, Cai. et al. decomposed EEG into five classical frequency bands and applied dynamic weighting to identify those most relevant to auditory attention tasks[52]. This method improved detection accuracy compared to models using fixed frequency bands. Similarly, Xie et al. developed a neural model that calculates the importance scores of these frequency bands using convolutional and fully connected layers, offering an efficient solution for real-time decoding with short time windows [54]. Other studies, such as that by Chen et al. [55], extracted DE features from the five standard bands through manual preprocessing and matched them with attention weights to improve classification performance. A generalized formulation of frequency attention mechanism can be expressed as follows. Given a frequency vector $X_{F}\in\mathbb{R}^{1\times F}$, the signal is divided into five frequency band subsets $F^{\delta},F^{\theta},F^{\alpha},F^{\beta},F^{\gamma}$, with each processed into a common feature dimension $f^{i}$, using a neural network or entropy function:
$f^{i}=function(F^{i})\qquad i\in\{\delta,\theta,\alpha,\beta,\gamma\}$ (4)
The resulting band-wise features are stacked into a matrix $X_{f}\in\mathbb{R}^{5\times f}$, followed by the initialization of an attention weight vector $W_{f}(f\in{\delta,\theta,\alpha,\beta,\gamma})$. The final frequency-aware feature representation is obtained via:
$\hat{X}_{f}=\sigma(X_{f}\odot W_{f})$ (5)
This process ensures that the model selectively emphasizes the most relevant spectral components, enhancing its ability to extract robust and discriminative features from EEG data.
Future developments of frequency attention mechanisms may focus on improving adaptability, efficiency, and interpretability. One promising direction is to move beyond static, predefined frequency bands by incorporating data-driven frequency decomposition methods that allow the model to discover meaningful spectral patterns tailored to specific individuals or tasks. Additionally, self-supervised learning techniques such as masked frequency modeling or contrastive learning could enable the model to extract useful frequency features from unlabeled EEG data, thus addressing the common challenge of limited annotations. To enhance interpretability, future work may also integrate generative models that visualize the relationship between attention weights and neurophysiological signals, providing clearer insights into the model’s decision-making process. These advancements are expected to improve the flexibility and transparency of EEG analysis, reinforcing the role of frequency attention as a critical component in advanced BCI systems.
@@ 2.2 Embedding Methods of Attention Modules
For example, embedding a channel attention module specifically enhances the model’s ability to capture important features at the channel level. Du et al. proposed the ATDD-LSTM model, which combined a channel attention module with a long short-term memory (LSTM) network. In this approach, the channel attention module was applied to feature vectors extracted by the LSTM layers, allowing the model to concentrate on channels most relevant to specific emotions while downplaying less relevant ones, which improved the accuracy of emotion recognition [56]. Although channel-level attention mechanisms in EEG analysis have shown promising results, conventional methods often flatten all channels into a uniform space when estimating their importance, disregarding the inherent spatial structure among electrodes, which may introduce bias. To overcome this limitation, Xu et al. integrated channel attention into a graph convolutional network (GCN), taking into account the spatial relationships between EEG recording electrodes [57]. Beyond channel attention, some studies focus on embedding temporal attention module to emphasize the temporal dynamics of EEG signals. Zhang et al. proposed a convolutional recurrent attention model that used CNNs to encode high-level representations and combined them with recurrent attention mechanisms (including LSTM networks and temporal attention modules). This method calculated attention weights on dynamic temporal features, allowed the model to focus on the most informative time periods, and extracted more valuable temporal information [58]. Traditional temporal attention modules typically rely on a purely data-driven learning paradigm, where the model autonomously identifies significant temporal segments. However, incorporating prior knowledge to guide the learning process can help the model identify key temporal patterns more efficiently. For example, inspired by the psychological peak-end rule, Kim et al. developed a model that integrated a bidirectional LSTM network with a temporal attention module. It assigned greater weight to emotional states occurring at key moments, capturing the dynamic variability of emotions over time and enhancing the model’s interpretability [59]. For frequency-domain features, frequency attention is rarely modeled in isolation. Instead, it is typically integrated with spatial and temporal features to enhance EEG signal representation.
@@
For example, Tao et al. proposed a deep learning model that incorporated both channel attention and inter-sample attention mechanisms. Since the samples were segmented based on time, the inter-sample attention effectively functioned as temporal attention. This approach enabled the model to effectively prioritize significant information across different channels and temporal segments for feature extraction [28]. Besides of the aforementioned method, Cai et al. introduced a dynamic attention mechanism that assigned different weights to different frequency sub-bands and channels of EEG signals [52]. This dynamic approach optimized feature representation and was applied within an adaptive decoding framework for complex downstream tasks. Additionally, to better capture the spatial relationships among EEG electrode channels, Jia et al. ’s GraphSleepNet [60] and Zhang et al. ’s hierarchical attention network based on graph structures [61] both utilized GCNs to model the spatial relationships of EEG electrodes. These approaches leveraged attention mechanisms across both time and space, significantly enhancing performance in tasks such as sleep stage classification and movement intention prediction. To extract information from EEG signals in a more comprehensive manner, a number of studies have implemented attention mechanisms across all three dimensions (channel, temporal and frequency). However, expanding attention to multiple dimensions requires careful architectural design to ensure a balanced and effective focus across them. Extending beyond channel and temporal information, Jia et al. introduced a spatio-temporal-spectral attention dense network that simultaneously considered temporal, frequency, and spatial features. This model adaptively captured crucial information across brain regions (spatial, i.e., channels), frequency bands (spectral), and time, providing a comprehensive feature extraction framework [49]. Xiao et al. extended Jia et al.’s model by proposing a neural network based on 4D attention [53]. In this approach, the channel dimension of the input samples is transformed into a two-dimensional feature to preserve the spatial positional information of the EEG signal electrodes, while also incorporating the time and frequency dimensions. It computed spatial attention (addressing the spatial positional relationships between channels) and frequency attention (applied to power spectral density and differential entropy features). These attention weights were then applied to refine the input, resulting in enhanced output features. Jia et al. ’s and Xiao et al. ’s models both leveraged temporal, frequency, and spatial characteristics of EEG channels, calculating attention weights across these dimensions to enhance the model’s focus on information pertinent to specific tasks. The main difference between their approaches lies in how they calculated attention weights and structured their models. Jia et al. computed attention separately across the frequency and time dimensions in parallel, and then merged these features for classification. In contrast, Xiao et al. integrated all dimensions into a unified 4D representation before computing attention, providing a different approach to capture feature interdependencies. While Jia et al. ’s dual-stream framework excels in modular feature extraction, it may miss cross-dimensional dependencies. In comparison, Xiao et al. ’s integrated 4D strategy enables holistic feature learning but introduces greater model complexity. The choice between these approaches depends on the dependency structure of the task: Jia et al. ’s method is more suited for problems where temporal and spectral features are relatively independent, whereas Xiao et al. ’s model is preferable when multi-dimensional interactions are crucial.
@@
A crucial aspect of applying attention mechanisms is defining the shape and dimensions of the input feature matrix. In previous research, attention weight computations have predominantly relied on the strict Euclidean geometric space of the feature matrix. However, given the brain’s complex topological structure, using Euclidean space alone may not accurately capture its underlying properties. Consequently, several studies have sought to align feature matrix definitions more precisely with the brain’s physiological structure by incorporating non-Euclidean space representations within attention mechanisms. For example, Zhang et al. introduced the concept of manifolds, proposing a time-frequency domain feature learning model that integrated both Riemannian manifold and Euclidean space representations. Their work demonstrated the effectiveness of attention mechanisms in synthesizing feature information across different mathematical domains [62].
@@
[61] 2019 Channel, Temporal CNN, GCN Movement decoding PhysioNet
[58] 2019 Temporal CNN, LSTM Movement intention recognition BCI IV-2A
@@
[60] 2020 Channel, Temporal GCN Sleep stage recognition MASS-SS
[28] 2020 Channel, Temporal RNN Emotion recognition DEAP, DREAMER
[59] 2020 Temporal LSTM, CNN Emotion recognition DEAP
[49] 2020 Channel, Temporal, Frequency CNN Emotion recognition SEED, SEEDIV
[62] 2020 Temporal, Frequency LSTM,CNN Emotion recognition SEED-VIG, SEED, BCI IV-2A, BCI IV-2B
[34] 2021 Channel CNN Auditory attention detection KUL
[52] 2021 Channel, Frequency CNN Auditory attention detection KUL, DTU
[27] 2021 Channel, Frequency CNN Motor imagery BCI IV-2A, BCI IV-2B, HGD
[53] 2022 Channel, Temporal, Frequency LSTM, CNN Emotion recognition DEAP, SEED, SEEDIV
[26] 2023 Channel, Temporal CNN Motor imagery BCI IV-2A, Custom
@@
[35] 2023 Channel CNN Motor imagery BCI IV-2A,BCI IV-2B, 2020BCI Track#1
[63] 2024 Channel CNN Motor imagery BCI IV-2A, BCI IV-2B, HGD, BCI III-4A
[64] 2024 Channel, Temporal CNN RSVP Tsinghua RSVP
[65] 2025 Channel CNN Motor imagery BCI IV-2A, BCI IV-2B, HGD
[66] 2025 Temporal CNN Motor imagery BCI IV-2A, BCI IV-2B